OpenStudy (gazela1997):

The zeros of parabola are 4 and -3 and the y-intercept is -10. determine the equation of a parabola using the factored form y=a(x-s)(x-t) HELLLPPP GOT AN MATH EXAM TOMORROW MEDAL AND FAN

OpenStudy (gazela1997):

@jim_thompson5910

OpenStudy (primeralph):

The zeros being at 4 and -3 means the equation is of the form y = a(x-4)(x+3). The intercept being at -10 means when x = 0, y = -10; so -10 = a(3)(-4) , so a = -10/12 = -5/6. So you have y = -5/6*(x+3)(x-4)